1) Nineteen coins are what’s in this cup.
Better believe me ‘cause that’s what’s up.
There’s no peeking or you’ll never learn.
Wait ‘til the end when it’s your turn.
A dollar thirty is the total inside.
That’s the truth, I’ve never lied.
So how many nickels and how many dimes?
Figure it out. I’ve got no more rhymes.

Answer by fractalier(6550)

Creative problem.
Let n and d be the number of nickels and dimes. Then we have
n + d = 19 and the value equation is
5n + 10d = 130 (we work in cents)
Now multiply the top equation by five and subtract from the second...we get
5n + 10d = 130
-(5n + 5d = 95)
-----------------
5d = 35
d = 7 dimes
n = 12 nickels